40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 272 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  3276   Fri Jul 23 14:26:01 2010 GopalUpdateOptic StacksSimple Frequency Response Measurements in COMSOL

Over the past couple days, I discovered a simple, direct method for calculating frequency responses with a combination of COMSOL and any plotter such as Excel or MatLab. The simple case of rectangular prism of steel was analyzed using this method; details will be posted shortly on the COMSOL Wiki page. The frequency response matched theoretical reasoning: the bar acts as a simple mechanical low-pass filter, rapidly attenuating driving frequencies at the base beyond the first eigenmode.

It therefore shouldn't be too difficult to extend this analysis to the MC1/MC3 stack. The many eigenfrequencies will produce a more complicated transfer function, and so more data points will be taken.

The major shortcoming of this method involves dealing with the imaginary components of the eigenfrequencies. As of now, I haven't found a way of measuring the phase lag between the drive and the response. I also haven't found a way of changing the damping constants and therefore playing with phase components.

 

  3275   Fri Jul 23 12:33:51 2010 josephbUpdateCDSmegatron, c1iscex, c1sus firewalled

Kiwamu and I strung an ethernet cable from the new 1X7 rack to the 1X3 rack. The cable is labeled c1iscex-daq on both ends.  This cable will eventually connect c1iscex's second ethernet port to the daq router.  However, for today, it plugged into the primary ethernet port and is going to a linksys router.  This is the same linksys router we used to firewall megatron.

The idea is to place megatron, c1sus, and c1iscex behind the firewall to prevent any problems with the currently running while doing RFM nework tests.

The way to get into the firewalled sub-network is to ssh into megatron.  The router will forward the ssh to megatron.  Inside the network, the computers will have the following IPs.  Router is 192.168.1.1, megatron is 192.168.1.2, c1sus is 192.168.1.3, and c1iscex is 192.168.1.4.

 

 

  3274   Fri Jul 23 10:16:44 2010 josephbUpdateCDSRerouted RFM around c1lsc, took RFM card out of c1lsc

 I re-routed around the c1lsc machine this morning.  I turned the crate off, and disconnected the transmission fiber from  c1lsc (which went to the receiver on c1asc).  I then took the receiving fiber from c1lsc and plugged it into the receiver on c1asc. 

I pulled out the c1lsc computer from the VME crate and pulled out the RFM card, which I needed for the CDS upgrade.  I then replaced the lsc card back in the crate and turned it back on.  Since there hasn't been a working version of the LSC code on linux1 since I overwrote it with the new CDS lsc code, this shouldn't have any significant impact on the interferometer.

I've confirmed that the RFM network seems to be in a good state (the only red lights on the RFM timing and status medm screen are LSC, ASC, and ETMX).  Fast channels can still be seen with dataviewer and fb40m appears to still be happy.

The RFM card has found its new home in the SUS IO Chassis.  The short fiber that used to go between c1asc and c1lsc is now on the top shelf of the new 1X3 rack.

  3273   Fri Jul 23 08:18:03 2010 josephbUpdateCDSNot enough room in IO chassis for RFM card - need to swap PMC to PCIe adapters

The End IO chassis have small trenton boards, which apparently only have 5 usuable PCI slots, even though there are 6 on the board.  This is because of the way the the host interface board is setup and its closeness to the 2nd to last PCI slot

The PMC to PCIe adapters I was handed by Jay for use with the RFM cards require a 4 pin power connection at the top, which are not available inside the thin 1U computers.

The only solution I can come up with is swap the PMC to PCIe adapters for the RFM cards with adapters for some of the already installed ADCs and DACs which do not require power directly from the power supply.  This should make it possible to mount the RFM card in the computer, at least for the ends.  Since the SUS and IOO chassis will have more slots available than needed, the RFM cards can be slotted into those.  The SUS has to fit in the chassis since the computer will have the Infiband host adapter and a dolphin connector for talking to the LSC machine.

There is still the problem of actually getting the RFM card into the computer, but that should be possible with a little bit of bending of the left side of the computer frame.

  3272   Fri Jul 23 08:15:59 2010 steveUpdatePSLref cavity ion pump

The ref cavity ion pump was running at 7.7kV instead of 5kV

This Digitel SPC-1 20 l/s ion pump should be running at 5kV

  3271   Fri Jul 23 00:13:11 2010 ranaUpdatePSLProblem NOT REALLY Solved

So...who was working around the PSL rack this morning and afternoon? Looks like there was some VCO phase noise work at the bottom of

the rack as well as some disconnecting of the Guralp cables from that rack. Who did which when and who needs to be punished?

  3270   Thu Jul 22 18:18:54 2010 AlbertoUpdatePSLProblem Solved

Quote:

Quote:

It looks like something wrong happened around the PSL front end.  One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy. 

We found it by the donkey alarm 10 minutes ago.

The attached picture is a screen shot of the PMC medm screen.

The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.

Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)

The problem seems to be a software one.

In any case, Kiwamu and I looked at the at the PMC crystal board and demod board, in search of a possible bad connection. We found a weak connection of the RG cable going into the PD input of the demod board. The cable was bent and almost broken.

I replaced the SMA connector of the cable with a new one that I soldered in situ. Then I made sure that the connection was good and didn't have any short due to the soldering.

[Alberto, Koji]

By looking at the reference pictures of the rack in the wiki, it turned out that the Sorensen which provides the 10V to the 1Y1 rack was on halt (red light on). It had been like that since 1.30pm today. It might have probably got disabled by a short somewhere or inadvertently by someone working nearby it.

Turning it off and on reset it. The crazy LO calibrated amplitude on the PMC screen got fixed.

Then it was again possible to lock PMC and FSS.

We also had to burtrestore the PSL computer becasue of the several reboots done on it today.

  3269   Thu Jul 22 15:59:29 2010 AlbertoUpdatePSLPSL front end machine

Quote:

It looks like something wrong happened around the PSL front end.  One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy. 

We found it by the donkey alarm 10 minutes ago.

The attached picture is a screen shot of the PMC medm screen.

The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.

Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)

The problem seems to be a software one.

In any case, Kiwamu and I looked at the at the PMC crystal board and demod board, in search of a possible bad connection. We found a weak connection of the RG cable going into the PD input of the demod board. The cable was bent and almost broken.

I replaced the SMA connector of the cable with a new one that I soldered in situ. Then I made sure that the connection was good and didn't have any short due to the soldering.

  3268   Thu Jul 22 14:07:20 2010 kiwamuUpdatePSLPSL front end machine

It looks like something wrong happened around the PSL front end.  One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy. 

We found it by the donkey alarm 10 minutes ago.

The attached picture is a screen shot of the PMC medm screen.

The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.

Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)

Attachment 1: PSL-PMC2010-07-22.png
PSL-PMC2010-07-22.png
  3267   Thu Jul 22 13:44:47 2010 JennaUpdatePEMGuralp seismometer

One of the Guralps [Gur2] has been taken to the atf gyro lab, along with the breakout box.

 

Edit by Jenne:  This means that we have no working seismometers in the 40m lab right now, so don't worry if you're looking for seismo data and you can't find any.  The 6 accelerometers should all still be up and running.

  3266   Thu Jul 22 08:07:27 2010 steveUpdateVACvent finished

Quote:

The vent has been finished successfully in this morning.

The vent was finished successfully this morning.

 Thanks to Kiwamu, Alberto and Koji

Attachment 1: 30hvent.jpg
30hvent.jpg
  3265   Thu Jul 22 07:19:56 2010 AlbertoUpdateTreasureMonsters, LNVR, and Phase noise

Quote:

On Picasa

 "They (shellfish) shall be an abomination to you; you shall not eat their flesh, but you shall regard their carcasses as an abomination." (Leviticus 11:11)

  3264   Thu Jul 22 03:08:27 2010 KojiUpdateIOOVent and MC lock

Summary

- The vacuum chambers have been vented.

- The north heavy door of the BS chamber has been opened by Genie (not by the crane).
It was replaced by the light door. The door is currently closed.

- The MC has been locked with 20mW incident and aligned. MC REFL was left unchanged but lock was able to be achieved.

- The optics before MCT CCD and MCT QPD have been adjusted for the low power operation.


Details

- The first HWP for the variable optical attanuator (HWP/PBS/HWP pair) was set to be 86deg from the maximum transmission at 126deg.
The incident power of 19mW has been measured.

- The PSL mechanical shutter has been manually opened. Two other beam blocks has been removed.

- I found the MC was totaly misaligned with no resonance.- I tried to align it based on the previous OSEM values but in vain.

How to align the MC mirrors from the scratch

- MC1 has been aligned so as to maximize REFL PD and DC signal of WFS QPDs.

- MC3 has been aligned by looking at the scattered light on the MC2 frames. The spot is centered on the MC2 approximately.

- MC2 has been aligned so that any resonance is seen in MC_F.

Modification of the MCT optics

- The ND filter before the MCT was removed.

- The Y1-45S mirror before the MCT CCD, which is also used to steer the beam to the MCT QPD path, was replaced to BS1-50-45P.
The reason I used 45P is to obtain higher reflectivity. Because S has higher reflectivity than P in the each layer, I expected to have higher reflectivity for S than 50%.

- The MC REFL path has not been untouched.

Modification of the servo

- The lock was attempted after alignment of the mirrors.  Here how to lock the MC is described below.

1. Run script/MC/mcloopson

2. Open the MC Servo screen in MEDM

3. Change the input gain from 6dB to 22dB.
Change offset from 0.78 to -0.464 (such that the length output has no offset).
Change VCO Gain from 3dB to 21dB

Change MC Length path Gain from 0.3 to 1.6

  3263   Thu Jul 22 01:02:08 2010 ranaUpdateTreasureMonsters, LNVR, and Phase noise

On Picasa

  3262   Wed Jul 21 19:11:18 2010 DmassUpdateGreen Lockinglocked

What did you use to filter the 2f components from your error signal? A homemade low pass or what?

 


Kiwamu:

I am using a homemade low pass filter.

It's just a RC passive LPF with the input impedance of 50 Ohm.

  3261   Wed Jul 21 17:41:17 2010 GopalConfigurationOptic StacksPictures of Stacks

Now that venting is complete, this is a request for anyone who opens any chamber:

1) Please notify me immediately so I can take pictures of the stacks in that chamber.

2) If I'm not around, please take a few pictures for me. I'm most interested in the shape, number of layers, size, and damper arrangements of each stack.

This is most important for the MC1/MC3 chamber, MC2 chamber, and BS/ITMX/ITMY chambers.

Thanks!

  3260   Wed Jul 21 15:43:38 2010 MeganSummaryPSLCopper Layer Thickness on the Reference Cavity

Using the equation for thermal resistance

Rthermal = L/(k*A)

where k is the thermal conductivity of a material, L is the length, and A is the surface area through which the heat passes, I could find the thermal resistance of the copper and stainless steel on the reference cavity. To reduce temperature gradients across the vacuum chamber, the thermal resistance of the copper must be the same or less than that of the stainless steel. Since the copper is directly on top of the stainless steel, the length and width will be the same for both, just the thickness will be different (for ease of calculation, I assumed flat, rectangular strips of the metal). Assuming we wish to have a thermal resistance of the copper n times less than that of the stainless steel, we have

RCu = RSS/n

or

L/(kCu*w*tCu) = L/(kSS*w*tSS*n)

so that

tCu/tSS = n*kSS/kCu

We know that kSS = 401 W/m*K and KCu = 16 W/m*K, so

tCu/tSS = 0.0399*n

By using the drawings for the short reference cavity vacuum chamber (the only one I could find drawings for online) I found a thickness of the walls of 0.12 in or 0.3048 cm. So for the same thermal resistance in both metals, the copper must be 0.0122 cm thick and for a thermal resistance 10 times less, it must be 0.122 cm thick. So we will have to keep wrapping the copper on the vacuum chamber!

  3259   Wed Jul 21 12:41:31 2010 kiwamuUpdateVACvent finished

The vent has been finished successfully in this morning.

The vent was finished successfully this morning.

  3258   Wed Jul 21 12:20:58 2010 RazibUpdatePhase CameraWeekly update

This past week I have worked on the following:

1. Setting up the infrastructure to do noise analysis: We added a temporary channel on the DAQ to connect to the PD 55 which we are using to take the power measurement. Before that, I connected the PD55 to an oscilloscope and recorded the power.

    phase_cam_setup_21_07_2010.jpg

The power at PD55 as measured using the oscilloscope = 600 µV.

Then I tried to calibrate the channel by sending up a signal from the function generator and measuring up the offset.. However, the channels seems noisy enough, especially due to electronics noise as suggested by the measurements and FFT calculation.

2. I worked on trying to sync the data acquisition of the PD and the CAM. After sometime spent on fiddling with the software method such as taking images at stamped time and then lining them up with the daq timestamps, I found a hardware method as suggested by Aidan. It was putting up a shutter (Uniblitz shutter and driver VMMD1) in the setup. I synced the shutter with the camera for which I had to tear apart the previously made trigger box and add a sync output from the camera (took a while as I also had to make a new cable).

3. I worked (still working) on making a differential amplifier to blow up the signal from the PD.

 

 

 

  3257   Wed Jul 21 12:20:29 2010 josephb, kiwamuUpdateComputersMegatron temporarily disconnected, c1iscex firewalled, green FE test

We are moving towards a first test of getting Kiwamu's green locking signals into the new front end at the new X end, as well as sending signal out to the green laser temperature control.

Towards that end, we borrowed the router which we were using as a firewall for megatron.   At the moment, megatron is not connected to the network.  The router (a linksys N wire router), was moved to the new X end, and setup to act as a firewall for the c1iscex machine.

At this point, we need to figure which channels of the DAC correspond to which outputs of the anti-imaging board (D000186) and coil driver outputs.  Ideally, we'd like to simply take a spare output from that board and bring it to the laser temperature control.  The watchdogs will be disabled when testing to avoid any unfortunate mis-sent signals to the coils.  It looks like it should be something like channels 6,7,8 are free, although I'm not positive if thats the correct mapping or if there's a n*8 + 6,7,8 mapping.

The ADC should be much easier to determine,  since we only have a single 16 channel set coming from the lemo breakout box.  Once we've determined channels, we should be all set to do a test with the green system.

  3256   Wed Jul 21 12:03:14 2010 Katharine, SharmilaUpdateWIKI-40M Updateweekly update

This past week, we levitated our small cylindrical magnet (with the flag made from heat shrink).  Though the levitated magnet didn't appear very jittery to the eye, we looked at the PD current on the scope and could see oscillations that corresponded to the flag hitting the sides of the OSEM.  The oscillations were more pronounced as we gently hit/vibrated the lab bench, and by pounding on the bench Rana knocked the levitated magnet completely out of the setup.  So, we're currently working on building a new, stabler mount.  The biggest challenge here is fixing the OSEM in place, but we're experimenting with different optics pieces to see which is the most stable for our purposes.   Jenne taught us how to make through holes using the drill press so we can add slats of aluminum to adjust the height of the OSEM mount.  We also plan to fix some heavy plates to the bottom of our system to decrease its vibration frequency.

We also calculated the transfer function of our circuit, which seems to match the measured frequency response to within a small factor.  We're playing with Rana's Simulink models and are currently modeling our own system to determine what gains we would expect to use and get a better understanding of our circuit.


Once our system is successfully mounted stably, we plan to experimentally observe the effects of changing the gain and integrator by looking at time series measurements of the PD current, as well as using the spectrum analyzer to compare the frequency response of our system at different gain settings.

  3255   Wed Jul 21 11:57:59 2010 GopalUpdateWIKI-40M Update7.14.10-7.21.10 Weekly Update

Summary of this week's activities:

7/14: Analytical calculation of Viton spring constant; updated Viton values in models; experimental confirmation of COMSOL eigenfrequencies (single stack layer)

7/15: Extensions to 2-, 3-, and 4-layer stack legs. Eigenfrequency characterizations performed for each level. Meshing issues with 4-layer stack prevented completion.

7/19: Debugged the 4-layer stack. Turned out to be a boundary condition issue because of non-sequential work-plane definitions. Successful characterization of single-leg eigenfrequencies.

7/20: Prototype three-legged stack completed, but dimensions are incorrect. Read Sievers paper for details of triple-legged stack. Sorted through many stack design binders in efforts to distinguish IOC/OOC, BSC/ITMX/ITMY, MC1/MC3, and MC2 dimensions.

7/21: Researched frequency domain analysis testing in COMSOL. Attempting to first find transfer function of a single-layer stack --> currently running into some run-time errors that will need some more debugging in the afternoon.

  3254   Tue Jul 20 23:52:36 2010 kiwamuUpdateVACupdate: slow vent has started

The vent is still going on. At this moment the pressure inside of the chambers is about 630 Torr.

Koji and I have replaced the 2nd instrument grade compressed air cylinder by the 3rd cylinder around 9 pm.

So far the vent speed has been nicely kept at about 1 Torr / min.

Attachment 1: untitled.png
untitled.png
  3253   Tue Jul 20 18:29:43 2010 nancyUpdateIOOQPD installed behind the MC2

 

Yesterday I installed teh QPD on the table behind MC2, and observed teh signal on it.

The MC_leak is directed to it by a steering mirror.

I used the A2L_MC2 script to minimise  teh pitch and yaw gains, and estimated teh spot position on teh MC2 using that.

This spot position was aligned to the center of teh QPD.

In the night while before taking measurements, I decided to turn off the Wavefront Sensor Servos, but just after that, the MC alignment went very bad, and I could not align it in the next 2 hours.

For some reason, the MC was really mad the whole day yesterday, and was getting misaligned again and again, even when the WFS feedback was on.

 

The table also had another IR laser in it, which I and Koji switched off.

 

I will continue measuring once we pump down again.

For now, I am analysing teh QPD circuit Transfer Function.

  3252   Tue Jul 20 17:38:16 2010 GopalConfigurationOptic StacksStack Type Clarifications

Some clarification is warranted regarding the different shapes of stacks. Corrections are appreciated:

1) The single-leg stack that I just completed should function as a working model for the IO, OO, and MC1/3. Rana commented, however, that the current dimensions are slightly off for MC1/3 (which makes sense since I could only find drawings for the IOC). If anyone knows the whereabouts of similar drawings for MC1/3, I'd much appreciate it.

2) A triple-leg stack can model the BS, ITMX, and ITMY chambers. I don't have exact dimensions for these, but I can make decent approximations from to-scale stack drawings. I'll probably work on a model for this style next, since at least I have some information regarding this version.

3) The MC2 chamber has its own stack model, about which I haven't found any drawings in the binders. I can't start on MC2C at all until I find such drawings.

  3251   Tue Jul 20 15:06:48 2010 josephb, bobOmnistructureGeneralNew speakers in ceiling in control room

After the crane training, Bob attached speakers to the ceiling right next to the projector, for use with presentations.

  3250   Tue Jul 20 11:55:15 2010 Sharmila,KatharineUpdateelogMaglev

We plotted the transfer functions for the maglev control circuit and compared them with the plots from the spectrum
analyzer. We were stuck for sometime because

1) we had wrongly entered the value of one of the resistors which was off by a factor of 2000.
2) The plots were not done in right units. So we couldn't figure out differences quite well.

The two plots are shown below. We are still off by a factor of 3 which we'll figure out soon.

P7140292.JPG

  3249   Tue Jul 20 11:49:31 2010 JenneUpdateSUSSome Suspensions not damping

[Jenne, Koji]

I moseyed into the control room this morning, to find ITMX and ITMY both with their watchdogs tripped.  ITMY (new convention) wouldn't damp.  Koji discovered that there was a sign flip in 2 of the sensors.  A set of reboots of c1susvme1&2 fixed the problem. 

A side note:  For the ETMs, the OSEM sensor readouts are gigantic (~20,000), whereas for the similar channels on all other optics, the readouts are on the order of 1.  After some looking around, it seems that this is just the way things have been (for at least 100 days), and the filters in the SUSPOS and other SUS filter banks have a high pass filter to take care of this.  It's weird, but it seems to be the way it is, and the ETMs damp, so it's all good.

  3248   Tue Jul 20 08:04:31 2010 steveConfigurationVACslow vent has started

The PSL shutter was closed. The beam path  blocked two places. High voltage power supplies to IOO and OMC PZT were checked to be off. Oplevs are off. 

The south arm green cavity was misaligned yesterday

We would like to keep the vent speed at 1 torr / min. I'm venting with N2 now up to 25 PSI. We have 3 cylinder  of instrument grade air inside the lab. Additional supply will arrive later. It can be as late as 1pm

  3247   Mon Jul 19 21:47:36 2010 ranaSummaryDAQDAQ timing test

Since we now have a good measurement of the phase noise of the Rb clock Marconi locked to the Rb clock, I wanted to use that to check out the old DAQ system:

I used Megan's phase noise setup - Marconi #2 is putting out 11000013 Hz at 13 dBm into the ZP-3MH mixer. Marconi #1 is putting out 3 dBm at 11000000 Hz into the RF input.

The output goes through a 50 Ohm load and then a Mini-Circuits BNC LP filter (either 2 or 5 MHz). Then an SR560 set for low noise, G = 5, AC coupling, 1-pole LP @ 1 kHz.

This SR560 output goes into the channel C1:IOO-MC_DRUM1 (which is sampled at 16384 Hz with ICS-110B after the usual Sander Liu AA chassis containing the INA134s).

  3246   Mon Jul 19 16:11:17 2010 GopalUpdateOptic StacksEigenfrequency Analysis of Full Stack

Expanding on the single-layer model, I added the second, third, and fourth layers to the stack in COMSOL. Eigenfrequency analysis run times increased exponentially as the model multiplied in complexity. The following images document the some of the important eigenfrequencies:


First Eigenmode: y-translational, 3.34 Hz:

Eigenfrequency_1_4Stacks.png


Second Eigenmode: x-translational, 3.39 Hz:

Eigenfrequency_2_4Stacks.png


Third Eigenmode: z-rotational, 3.88 Hz:

 Eigenfrequency_3_4Stacks.png


Sixth Eigenmode: z-translational, 8.55 Hz:

Eigenfrequency_6_4Stacks.png


As expected, the eigenfrequencies are generally lower, but still in the same range, as the single-layer model, because of greater mass but constant weight-per-spring distribution.

Next Steps:

1) Extend a single stack to the full stack system, which consists of three stacks like this. Perform similar eigenmode analysis.

2) Analyze the mirror suspension system and incorporate a similar pendulum on the top plate.

3) Make transfer function measurements between seismic and mirror motions.

  3245   Mon Jul 19 14:16:01 2010 nancy, kojiUpdateIOOQPD Response Transfer Function

Quote:

Friday night myself and Koji measured the Transfer function of the QPD circuit at MC2 side using a chopper . Following was our procedure :

 

We connected some wires at the input and output of teh filter circuit to one of the segment of teh QPD. - seg 2.

A laser light was shined on to  the QPD, it was pulsed using a chopper. The frequency of rotation of teh chopper was varied.

These wires were then fed to the spectum analyser , and a transfer funstion was observed, It was nearly a low pass filter

The chopper frequency was then made variable by giving the chopper a signal from teh spectrum analyser. This signal just swiped a large range of the rpm of the chopper.

Now the input signal looked like a sine wave of varying frequency. the transfer functino looked like a perfect LPF, with a small SNR.

Attaching the plot of the TF in the next e-log (this one is on windows and cant access /cvs/cds)

 

 QPDTF2.png

  3244   Mon Jul 19 14:14:03 2010 nancy, kojiUpdateIOOQPD Response Transfer Function

Friday night myself and Koji measured the Transfer function of the QPD circuit at MC2 side using a chopper . Following was our procedure :

 

We connected some wires at the input and output of the filter circuit to one of the segment of teh QPD. - seg 1.

A laser light was shined on to the QPD, it was pulsed using a chopper. The frequency of rotation of the chopper was varied.

These wires were then fed to the spectum analyser , and a transfer funstion was observed, It was nearly a low pass filter

The chopper frequency was then made variable by giving the chopper a signal from the spectrum analyser. This signal just swiped a large range of the rpm of the chopper.

Now the input signal looked like a sine wave of varying frequency. the transfer function looked like a perfect LPF, with a small SNR.

Attaching the plot of the TF in the next e-log (this one is on windows and can't access /cvs/cds)

 

  3243   Mon Jul 19 13:51:09 2010 kiwamuUpdateCDSTIming card at X end

[Joe, Kiwamu]

 

 The timing slave in the IO chassis on the new X end was not working with symptoms of no front "OK" green light, no "PPS" light, 3.3V testpoint not working and  ERROR testpoint bouncing between 5-6V.

We took out the timing slave from the X end IO chassis put in to the new Y end IO chassis .

It worked perfectly there. We took the working one from Y end put in the X end IO chassis.

We slowly added cables. First we added power , it worked fine and we saw green "OK" light. Then we added 1PPS signal by a fiber and it also worked.

We turned everything off and then we added 40pin IPC cable from the chassis and infiniband cable from the  computer.

When we turned ON it we didn't see the green light.

This means something in the computer configuration might be wrong not in the timing card, we now are trying to make contact with Alex.

We are comparing the setup of the C1SCX  machine and the working C1ISCEX machine.

  3242   Sun Jul 18 20:50:03 2010 ranaConfigurationIOOMC TRANS optics changed

To make the beam on the MC trans camera bigger, I removed the lens + ND filter that was in that path.

The camera was getting the transmission through a BS1-33 (33% reflector). The reflection went to the TRANS QPD. I changed

the R=33% into an HR mirror (Y1-45P) so now the camera has a nice beam. The QPD was now saturating so I put a ND06 into that path

so now the TRANS_SUM is ~4.5-5 V when the MC is aligned.

The MC was also misaligned and failing to lock all weekend (why??) so I aligned the MC mirrors to get it to acquire again. Since we want to

collect MC seismic data, please make sure the MC is locked and running after finishing with your various MC or PSL work (this means YOU).

  3241   Fri Jul 16 23:53:27 2010 RanaUpdatePSLReference Cavity Insulation

From the trend, it seems that the Reference Cavity's temperature servo is working fine with the new copper foil. I was unable to find the insulating foam anywhere, but that's OK. We'll just get Frank to make us a new insulation with his special yellow stuff.

The copper foil that Steve got is just the right thickness for making it easy to form around the vacuum can, but we just have to have the patience to wrap ~5-10 more layers on there. We also have to get a new heater jacket; this one barely fits around the outside of the copper wrap. The one we have now seems to have a good heating wire pattern, but I don't know where we can buy these.

I also turned the HEPA's Variac back down to the nominal value of 20. Please remember to turn it back up to 100 before working on the PSL.

  3240   Fri Jul 16 20:25:52 2010 MeganUpdatePSLReference Cavity Insulation

Rana and I

1) took the temperature sensors off the reference cavity;

2) wrapped copper foil around the cavity (during which I learned it is REALLY easy to cut hands with the foil);

3) wrapped electrical tape around the power terminals of the temperature sensors (color-coded, too! Red for the out of loop sensor, Blue for the first one, Brown for the second, Gray for the third, and Violet for the fourth. Yes, we went with an alphabetical coding system, excluding the out of loop sensor);

4) re-wrapped the thermal blanket heater;

5) covered the ends of the cavities with copper, ensuring that the beam can enter and exit;

6) took pretty pictures for your enjoyment!

We will see if this helps the temperature stabilization of the reference cavity.

 

DSC_2271.JPG

The end of the reference cavity, with a lovely square around the beam.

 

DSC_2266.JPG

The entire, well-wrapped reference cavity!

  3239   Fri Jul 16 16:12:31 2010 AlbertoConfigurationComputersc1susvme1/2 rebooted

Today I noticed that the FE SYNC counters of c1susvme1/2 on the RFM network screen were stuck at 16384. I tried to reboot the machines to fix the problem but it didn't work.

The BS watchdog tripped off when I did that, because I had forgotten to disable it. I had to wait for a few minutes before it settled down again.

Later I also re-locked the mode cleaner. But before I could do it, Rana had to reduce the MC_L offset for me.

  3238   Fri Jul 16 16:07:14 2010 josephbUpdateComputersPossible solution for the last ADC

After talking with Jenne, I realized the ADC card in the c1ass machine was currently going unused.  As we are short an ADC card, a possible solution is to press that card into service.  Unfortunately, its currently on a PMC to PCI adapter, rather than PMC to PCIe adapter.  The two options I have are to try to find a different adapter board (I was handed 3 for RFM cards, so its possible there's another spare over in downs - unfortunately I missed Jay when I went over at 2:30 to check).  The other option is put it directly into a computer, the only option being megatron, as the other machines don't have full length PCI slot. 

I'm still waiting to hear back from Alex (who is in Germany for the next 10 days) whether I can connect both in the computer as well as with the IO chassis.

So to that end, I briefly turned off the c1ass machine, and pulled the card.  I then turned it back on, restarted all the code as per the wiki instructions, and had Jenne go over how it looked with me, to make sure everything was ok.

There is something odd with some of the channels reading 1e20 from the RFM network.  I believe this is related to those particular channels not being refreshed by their source (which is other suspension front end machines), so its just sitting at a default until the channel value actually changes.

 

 

  3237   Fri Jul 16 15:57:19 2010 josephb,kiwamuUpdateComputersNew X end FE and IO chassis work

We finished setting up the new X end front end machine (still temporarily called c1scx), and attached it to its IO chassis.  We're preparing for a test tomorrow, where we redirect the Limo breakout box to the new front end and IO chassis, so Kiwamu can test getting some green locking channels into his controls model.

We strung a pair of blue fibers from the timing master to the new X end (and labeled them), so we have a timing signal for the IO chassis.  I also labeled the orange fiber Alex had repurposed from the RFM to timing for the new Y end when I noticed he had not actually labelled it at the timing master.

  3236   Fri Jul 16 15:39:27 2010 nancyUpdateIOOWavefront Sensors- switched off

I tuned the gain of WFS to 0 last night at about 3am.

I turned it back on now.

  3235   Fri Jul 16 13:05:48 2010 Kyung-haUpdateSUSLate update for 7/13 Tue (Tip Tilts)

[Jenne & Kyung-ha]

We suspended the mirror to one of the main frame with the ECD backplane we finished before. The hard task was to find the right balance for the mirror so that 1) it won't be tilted and 2) it'll be in the right position for the ECD backplanes so that the magnets attached to the mirror holder would be in the very center of each ECD holes. We used optical lever laser (red He/Ne) to check the balance of the mirror. We tried to use the jig for the mirror holder clamps but because of the size difference, we couldn't use it at all. (Since the magnets are very heavy, we thought the wire being not perfectly centered might work better. However, the jig dimension was way too different that the wire ended up in the middle of one of the holes.) Since there was no other clever way to attach the wire in the right position, we just tried to be as center/accurate as possible. After attaching wire to that mirror holder clamps, we hanged it to the frame. Again, we couldn't find any other accurate way to find the center so we held the wire and tried to adjust the mirror height as accurate as possible so that it can be in the right position in respect to ECD backplane and not be tilted at the same time. However, when we hanged the mirror, it was still tilted.. So we adjusted the mirror tilt using the mirror holder clamps. Since the holes on the clamps were ellipse shapes, we could adjust the position of the clamps a little bit. When we adjust the clamps, we started to tighten the screws when the mirror is NOT in the perfect position since the tightening up part changes the mirror angle anyways. Luckily, when we tightened up the last screw, the mirror was in the perfect position! After that, we poked the mirror several times to make sure that it comes back to the same place.

Amazingly, we could finish this whole hanging/adjusting process in about 30 mins! :D (Jan said it's because of his amazing moral support. :P Maybe he'll be there to support us everytime we work on the mirrors?)

  3234   Fri Jul 16 12:36:00 2010 Katharine, SharmilaUpdateeloglevitation

After last night's challenge (or inspiration), we levitated our magnet this morning.  Since the nice Olympus camera is not currently in the 40m, we had to use my less stellar camera, but despite the poor video quality you can still see the magnet returning to its stable equilibrium position.  Once we recover the better camera, we will post new videos.  Also, we haven't yet figured out how to put videos in line in the elog entry, so here are the youtube links:

 

levitation 1

levitation 2

 

We adjusted the gain on coil 1 so that the resistance from the pots was 57.1k (maximum gain of 101.2,).

currents from power supply, pre-levitation: 0.08 A and 0.34 A

post levitation: 0.08 A and 0.11 A


note: we're not sure why changing the gain on coil 3 changes the current through the power supply, so we'd like to investigate that next.

Attachment 1: CIMG0649.AVI
  3233   Thu Jul 15 23:51:47 2010 Mr. MaricHowToSUSLevitate me if you can

You guys must work harder.

mag_lev.jpg


  3232   Thu Jul 15 19:27:04 2010 ranaSummaryPSLRC trend after the insulation removal

Untitled.png

As you can see, there was not much (if any) worsening of the laser frequency fluctuation from removing the RefCav insulation. The plots below are zooomed in:

a.pngb.png

I have used the same peak-to-peak scale so that its easy to compare the fluctuations before (LEFT) and after (RIGHT).

As you can clearly see, the laser frequency moves just as much now (the SLOW_DC) as it did before when it had the insulation. Only now the apparent (i.e. fake) RC temperature fluctuations are much larger. So this sensor is fairly useless as configured.

  3231   Thu Jul 15 19:13:03 2010 ranaUpdateIOOMC_F check

Sometimes I like to plot the spectrum of MC_F. Its a good diagnosis of whether something is wrong.

The red trace is noisier than the blue reference. What is the cause of this?

Attachment 1: a.png
a.png
  3230   Thu Jul 15 16:57:31 2010 josephb, kiwamuUpdateCDSThe temporarily named c1scx machine getting ready to connect to IO chassis

This is the machine which will be the new x end front end machine.  Its IP is 192.168.113.86. 

We changed the root and controls passwords to the usual.  We have modified the controls user group to be 1001, by using "usermod -u 1001 controls" (we had to use the non-rtl kernel to get that command to work).

We changed /etc/fstab to point to /cvs/cds on Linux rather than some downs machine.  We added a link to /cvs/cds/rtcds in the local /opt directory.

We modified the /etc/rc.d/rc.local file to no longer run /opt/open-mx/sbin/omx_init start, /cvs/cds/geo/target/fb/mx_stream -d scipe12:0 om1, and /cvs/cds/geo/target/fb/mx_stream -d scipe12:0 -e2 -r2 om2.  We modified the /usr/bin/setup_shmem.rtl to run only c1x00 c1scx and c1spx.

I also commented out a line0 "/bin/rm -f /rtl*"

 

  3229   Thu Jul 15 16:16:51 2010 AlbertoConfigurationLSCPRC and SRC length adjustements

Quote:

Tell me whether it is correct or not. Otherwise I won't be able to sleep tonight.

Quote:

According to these results, these would be the proposed adjustements to the cavity lengths:
dl(PRC) = -0.0266 m; dl(SRC) = 0.612 m

 Sorry. I was in a rush to go to the LIGO "all hands" meetings when I posted that elog entry, that I forgot a zero in the SRC length value. The correct values are:

dl(PRC) = -0.0266 m; dl(SRC) = 0.0612 m

The cavity absolute lengths are then:

L(PRC) = 0.5/2/f1*c - 0.0266 = 6.7466 m

L(SRC) = c/f2 + 0.0612 = 5.4798 m

where c is the speed of light; f1 = 11065399 Hz; f2 = 55326995 Hz

  3228   Thu Jul 15 15:57:10 2010 KojiConfigurationLSCPRC and SRC length adjustements

Tell me whether it is correct or not. Otherwise I won't be able to sleep tonight.

Quote:

According to these results, these would be the proposed adjustements to the cavity lengths:
dl(PRC) = -0.0266 m; dl(SRC) = 0.612 m

  3227   Thu Jul 15 12:21:08 2010 AlbertoConfigurationLSCPRC and SRC length adjustements

Lately I've been trying to calculate the corrections to the recycling cavity lengths that would compensate for the phase that the sidebands will pick up from the arms in the upgraded interferometer.

To do that calculation , I tried two quite different ways, although equivalent in principle. They both use the optickle model of the 40m, but the calculation is made differently.

In the first way, I looked directly at the phases of the field: phase of [input field] / [reflected field], phase of [input field at PRM] / [transmitted field at SRM].

In the second way I looked at the demodulation phases of the LSC signals.

The first way is much simpler, especially from a computational point of view. It is the first I tried several weeks ago, but then I had abandoned because back then I thought it wasn't the correct way.

Anyway, both ways gave me the same results for the PRC length.
For the SRC length, the first way has given me a clear outcome. On the other hand, the second way has produced a less clear result.

According to these results, these would be the proposed adjustements to the cavity lengths:
dl(PRC) = -0.0266 m; dl(SRC) = 0.0612 m

I) 1st Way
a) case of arms ideal length (33.86 m)

sidebandPhaseRotation_73433447384.png sidebandPhaseRotation_73433447590.png


b) case arm length = 38.40 m

PRC

 sidebandPhaseRotation_73433447960.png zoom -> sidebandPhaseRotation_73433448864.png

SRC

sidebandPhaseRotation_73433445372.png zoom -> sidebandPhaseRotation_73433449354.png


II) 2nd Way
a) case of arms ideal length (33.86 m)

 demodPhaseVsRecCavLengths_73432637363.png


b) case arm length = 38.40 m

demodPhaseVsRecCavLengths_73433237349.png

ELOG V3.1.3-